# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@
Abstract.
Visually Rich Documents (VRDs) are essential in academia, finance, medical fields, and marketing due to their multimodal information content. Traditional methods for extracting information from VRDs depend on expert knowledge and manual labor, making them costly and inefficient. The advent of deep learning has revolutionized this process, introducing models that leverage multimodal information—vision, text, and layout—along with pretraining tasks to develop comprehensive document representations. These models have achieved state-of-the-art performance across various downstream tasks, significantly enhancing the efficiency and accuracy of information extraction from VRDs. In response to the growing demands and rapid developments in Visually Rich Document Understanding (VRDU), this paper provides a comprehensive review of deep learning-based VRDU frameworks. We systematically survey and analyze existing methods and benchmark datasets, categorizing them based on adopted strategies and downstream tasks. Furthermore, we compare different techniques used in VRDU models, focusing on feature representation and fusion, model architecture, and pretraining methods, while highlighting their strengths, limitations, and appropriate scenarios. Finally, we identify emerging trends and challenges in VRDU, offering insights into future research directions and practical applications. This survey aims to provide a thorough understanding of VRDU advancements, benefiting both academic and industrial sectors.
Visually Rich Document Understanding, Key Information Extraction, Question Answering, Entity Linking, Multimodal
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Information extraction
1. Introduction
1.1. Backgrounds
Visually Rich Documents (VRDs) are documents that contain a diverse mix of visual and textual elements designed to convey information in a comprehensive and visually engaging manner, which are ubiquitous in our daily lives and appear in domains such as finance, medicine, and academia. These documents integrate various types of visually rich elements, including paragraphs, tables, charts, diagrams, and photos. Both textual elements (such as paragraphs, lists, and captions) and visually rich elements (such as tables and figures), collectively referred to as document semantic entities, are essential for illustrating, explaining, and summarizing information. Common formats for VRDs include PDF (Portable Document Format), DOC/DOCX (Microsoft Word Document), and image files (JPEG, PNG). These documents are typically structured in a semi-structured or even unstructured manner, making their understanding and information extraction challenging.
The VRD Understanding (VRDU) field is dedicated to comprehending the structure of VRDs and extracting pertinent information, thus transforming unstructured or semi-structured content into a machine-readable format. VRDU mainly encompasses two tasks: Key Information Extraction (KIE) and Question Answering (QA). Key Information Extraction (KIE) aims to identify and extract values based on predefined keys. Depending on the model used, KIE can be approached as an entity retrieval task (KIE Task Formulation 1) or a sequence tagging task (Task Formulation 2), as shown in Figure 1. Visually Rich Document Question Answering (VRD-QA) involves answering questions in natural language based on contextual information in VRDs. As explained in Figure 1, the model must locate the answer within the document to respond to the input question. Beyond extractive methods, KIE and VRD-QA can also be formulated as generative tasks, where the model generates the desired value or answer auto-regressively based on the query or question and the document images. In addition to the above-mentioned tasks, another content understanding task is Entity Linking (EL), which is proposed to identify the semantic correlations between document entities (parent-child relations).
To address visually rich document understanding tasks, heuristic methods (Watanabe et al., 1995; Seki et al., 2007; Rusinol et al., 2013) and statistical machine learning techniques (Oliveira and Viana, 2017) have been used in domain-specific document applications, which require expert customization and are challenging to update. Recent advances in deep learning have introduced promising alternatives. LSTM and CNN-based models (Katti et al., 2018; Denk and Reisswig, 2019; Zhao et al., 2019), feature-driven approaches (Yu et al., 2021; Zhang et al., 2020; Wang et al., 2021a), and layout-aware pretrained frameworks (Xu et al., 2020a; Wang et al., 2022b; Hong et al., 2021), along with visual-integrated pretrained frameworks (Huang et al., 2022), have significantly improved document representation and achieved state-of-the-art performance in various downstream tasks. However, these models primarily focus on fine-grained features, such as grids and word/word pieces (textual tokens), and often struggle with computational complexity and the ability to capture global logical and layout correlations. In contrast, coarse-grained models, which operate at the entity level, address some limitations of fine-grained models but may omit detailed information, resulting in less comprehensive representations. To bridge these gaps, joint-grained frameworks (Li et al., 2021c; Yu et al., 2022; Bai et al., 2022) and LLM-based frameworks (Luo et al., 2024; Liu et al., 2024b) have been developed.
With the rapid development and increasing demands in Visually Rich Document Understanding (VRDU), various model architectures, multimodal learning methods, and pretraining techniques have been introduced to consistently enhance the performance of specific or multiple VRDU tasks. A comprehensive review of VRDU frameworks based on deep learning is provided, systematically surveying and analyzing existing methods and benchmark datasets, which are categorized based on adopted strategies and downstream tasks. Additionally, different techniques used in VRDU models are compared, focusing on feature representation and fusion, model architecture, and pretraining methods, with their strengths, limitations, and appropriate scenarios highlighted. Finally, emerging trends and challenges in visually rich document content understanding are identified, offering insights into future research directions and practical applications.
1.2. Scope
This paper is devoted to surveying and analysing existing work on understanding deep learning-based document content. Papers on the following topics will be reviewed and summarized in this paper.
  • The paper focuses on document content understanding tasks that will be included in this paper, mainly including document key information extraction, visual question answering and entity linking. This paper will not summarise the models and datasets that only focus on document structure understanding tasks.
  • A paper proposed a novel deep-learning VRDU model for addressing one or more Visually Rich Document Content Understanding (VRD-CU) tasks. Section 2.2.1 will briefly describe typical document content understanding approaches in a heuristically or traditional machine learning way.
  • This paper compiles widely used VRDU benchmark datasets and recently proposed VRDU datasets from top-tier conferences and journals since 2019. All included datasets are publicly available.
@@ 1.2. Scope
  • This paper focuses on the multimodal capabilities for achieving a holistic understanding of entire, multi-page documents, in contrast to previous document understanding surveys that primarily focus on isolated subtasks.
@@
It should be noted that this article only investigates models and datasets designed for the entire document or the document page rather than the specific document components such as table, chart. Therefore, models and datasets proposed for table detection, table structure recognition, and chart or plot question answering will not be summarized in the main body of this survey.
1.3. Related Surveys
@@
Several surveys offer a comprehensive overview of general document understanding tasks. These surveys primarily focus on document layout analysis (Subramani et al., 2020), table extraction (Liu et al., 2023), named entity recognition (Ehrmann et al., 2023), and document image analysis (Lombardi and Marinai, 2020) across diverse document types, including invoices (Saout et al., 2024) and historical document (Ehrmann et al., 2023; Lombardi and Marinai, 2020). In particular, computer vision-based research focuses on scanned document analysis and structure understanding. While these studies have advanced document image analysis, they often focus on fragmented subtasks and fall short of providing a holistic understanding of an entire document on multiple pages.
@@
Recent advancements in deep learning have fueled the emergence of VRDU tasks, which demand complex document content understanding capabilities. These tasks include key information extraction, question answering, and document entity linking. However, existing surveys have not adequately addressed the unique challenges and opportunities presented by VRDU, with a specific emphasis on deep learning-based multimodal approaches (Cui et al., 2021). To bridge this gap, this survey aims to provide a comprehensive overview of VRDU frameworks and datasets, including multimodal feature extractions and fusions in both mono and multi-task VRD models.
1.4. Contributions
@@
  • The paper provides a detailed review and systematic categorization of VRDU frameworks and benchmark datasets, organized based on adopted strategies and downstream tasks.
  • It critically examines and compares different techniques used in VRDU models, focusing on feature representation and fusion, model architecture, and pretraining methods, highlighting their strengths, limitations, and appropriate scenarios.
@@
1.5. Survey Structure
Section 1 in this survey discusses the background of visually rich document understanding and emphasizes the aim, scope, and contributions of this paper. Section 2 provides additional background knowledge, including definitions of the VRDU tasks covered in this research and the development of document understanding. Sections 3 and 3 review significant frameworks for mono-task and multi-task VRDU, respectively. Section 5 compiles benchmark datasets for three VRDU subtasks: scanned receipts, forms, and single and multiple-page documents from diverse domains. Section 6 critically examines and compares the techniques adopted by various models, highlighting their strengths and limitations. Finally, Section 7 summarizes the trends and future directions in this field.
2. Background
2.1. Task Definition
Based on the differences in the purpose and application scenarios, visually rich document understanding tasks are classified into three categories, including Key Information Extraction, Question Answering and Entity Linking.
Figure 1. Visually rich document content understanding task clarifications.
@@ 2.1. Task Definition
  • Question Answering: is a task of answering questions about a VRD by using natural languages. Based on the answer types, it can be divided into extractive VQA and abstractive VQA. The answer from extractive VQA is directly extracted from the target document, while abstractive VQA requires generative answers based on comprehensively understanding questions and related VRDs.
@@
  • Document Entity Linking: refers to identifying the semantic relations between document entities to construct the logical structure of the input document image.
2.2. Development of Document Understanding
2.2.1. Traditional Approaches
@@
Rule-based methods, as highlighted in several studies (Watanabe et al., 1995; O’Gorman, 1993; Seki et al., 2007; Rusinol et al., 2013), have demonstrated high precision in domain-specific applications. However, these methods have several drawbacks: they are manually intensive, costly, and require expert intervention for tailored customization. Additionally, they are inflexible, often necessitating frequent manual updates, even for minor modifications. In response to these limitations, machine learning-based approaches have been proposed for document understanding. For instance, SVM-based methods have been utilized for layout understanding (Oliveira and Viana, 2017), and TF-IDF techniques combined with hand-crafted features have been applied for extracting information from invoices. Furthermore, rule-based and statistical models have been used for entity extraction. Despite these advancements, machine learning methods rely heavily on human intervention and domain-specific expertise. They are time-consuming and often deliver suboptimal performance. Moreover, the majority of these methods typically depend on single-modality data, restricting them to either layout, text, or visual inputs.
2.2.2. Single Modality-based Approaches
@@
With the advancement of deep learning, deeper model architectures such as CNNs (Katti et al., 2018; Yang et al., 2017) have emerged, and pretrained language (Devlin, 2018; Liu et al., 2019b) and vision models (Ren et al., 2015; He et al., 2017) are now commonly used as robust baselines for understanding VRD content and structure. Due to the multimodal nature of VRDs, which involves the integration of text, vision, and layout, researchers are increasingly focusing on leveraging this combined information to achieve significant improvements in various downstream VRDU tasks.
2.2.3. Cross Modality based Approaches
@@
Considering the multimodal nature of VRDs, many frameworks propose various methods to encode multimodal information, including text, vision, and layout, and fuse them effectively. Text and vision information is normally encoded by pretrained backbones such as BERT (Devlin, 2018) or RoBERTa (Liu et al., 2019b) for textual features, Faster-RCNN (Ren et al., 2015) or Mask-RCNN (He et al., 2017) for visual features. For layout information, different encoding methods are introduced, including linear projection (Wang et al., 2020b), 2D positional encoding (Xu et al., 2020a), and attention bias to allow the proposed models to be layout-sensitive. Different feature fusion methods are introduced including summing up (Yu et al., 2021), concatenation (Lee et al., 2022b), attention-based contextual learning (Majumder et al., 2020a), and prompting (He et al., 2023). However, most of those frameworks leverage implicit knowledge from pretrained backbones with a task-orientated shadow adapter for specific VRDU downstream tasks such as KIE (Lee et al., 2023; Wang and Shang, 2022; Chen et al., 2023; Cao et al., 2023b) or EL (Zhang et al., 2021; Hu et al., 2023; Carbonell et al., 2021). Those frameworks tend to achieve delicate performance on specific tasks or document formats instead of acquiring a generalised model to represent documents comprehensively.
2.2.4. Multimodal Pre-training Approches
@@
Inspired by the success of BERT-style models (Devlin, 2018; Liu et al., 2019b) in acquiring knowledge through self-supervised learning, pretrained document understanding models have emerged to harness self-supervised or supervised pretraining tasks from extensive document collections. LayoutLM (Xu et al., 2020a), the first encoder-only VRDU model, utilizes self-supervised tasks, such as masked vision-language modelling, with text and layout information. Subsequent models have expanded on this by integrating layout information (Wang et al., 2022b; Tu et al., 2023; Li et al., 2021c) and visual cues (Xu et al., 2020b; Huang et al., 2022) through multimodal transformers. While encoder-only models have shown significant improvements on various benchmark datasets (Jaume et al., 2019; Park et al., 2019; Huang et al., 2019; Mathew et al., 2021; Harley et al., 2015), they often require detailed annotations and are limited by fixed input lengths. To address these limitations, encoder-decoder frameworks (Tang et al., 2022; Kim et al., 2022; Davis et al., 2022) and prompt-based methods for LLMs/MLLMs (Luo et al., 2024; Liu et al., 2024b) have been developed, enhancing layout awareness and performance in VRDU tasks. However, a significant gap remains in effectively applying these models in real-world scenarios with zero shot.
3. Mono-Task Document Understanding Frameworks
Visually rich document content understanding encompasses several independent downstream tasks tailored to different application scenarios and user demands. This section introduces methods focused on specific document understanding tasks. Many of these models aim to provide comprehensive document representation and integrate target-oriented modules or techniques to improve performance and efficiency across various VRDU tasks. The models for three VRDU downstream tasks—Key Information Extraction (KIE), Entity Linking (EL), and Visual Question Answering (VQA)—are introduced and summarized with insights into current trends.

### v2

@@
Abstract
Visually Rich Documents (VRDs) play a vital role in domains such as academia, finance, healthcare, and marketing, as they convey information through a combination of text, layout, and visual elements. Traditional approaches to extracting information from VRDs rely heavily on expert knowledge and manual annotation, making them labor-intensive and inefficient. Recent advances in deep learning have transformed this landscape by enabling multimodal models that integrate vision, language, and layout features through pretraining, significantly improving information extraction performance. This survey presents a comprehensive overview of deep learning-based frameworks for VRD Content Understanding (VRD-CU). We categorize existing methods based on their modeling strategies and downstream tasks, and provide a comparative analysis of key components, including feature representation, fusion techniques, model architectures, and pretraining objectives. Additionally, we highlight the strengths and limitations of each approach and discuss their suitability for different applications. The paper concludes with a discussion of current challenges and emerging trends, offering guidance for future research and practical deployment in real-world scenarios.
keywords:
Visually Rich Document, Deep Learning, Multimodal Learning
1 Introduction
1.1 Backgrounds
Visually Rich Documents (VRDs) are widely used across various domains, including finance, healthcare, law, education, and research. Both structured and unstructured visual-textual information must be interpreted in combination to derive meaningful insights. Unlike plain-text documents, VRDs contain heterogeneous components such as paragraphs, tables, charts, diagrams, and images, which are intricately arranged in formats like PDFs [mmvqa, pdfvqa], Word documents [docbank], or scanned images [vrdiu]. These visual and textual elements, referred to as document semantic entities, play a critical role in summarizing, illustrating, or substantiating key content, often through complex layouts and visual cues.
Understanding VRDs presents significant challenges due to their semi-structured or unstructured nature, which limits the effectiveness of traditional rule-based or linear text processing methods. This complexity has given rise to Visually Rich Document Content Understanding (VRD-CU), a specialized field focused on automatically extracting, interpreting, and reasoning over such documents to enable machine-readability and scalable knowledge extraction.
Historically, early VRD-CU approaches were heuristic-driven, relying on layout analysis and domain-specific templates [watanabe1995layout, seki2007information], later expanding to include statistical models [rusinol2013field, oliveira2017fast]. However, these methods required manually crafted rules, lacked flexibility, and often failed to generalize beyond fixed document structures or domains. The emergence of deep learning marked a turning point in VRD-CU. Pioneering models such as Chargrid [chargrid] and BERTgrid [bertgrid] applied convolutional and recurrent neural networks to treat document layouts as two-dimensional text grids. This led to the development of feature-based models such as PICK [pick], TRIE [trie], and Ephoie [ephoie], which integrated layout features with textual signals for improved extraction.
To acquire comprehensive VRD representations and improve generalization across VRD tasks, BERT-style pretrained frameworks have been proposed to achieve comprehensive document representation in VRD. They align multimodal information through various pretraining tasks. However, their focus on fine-grained word sequences limits the ability to capture higher-level semantic relationships. To address these limitations, coarse-grained models are introduced[udoc, selfdoc] that focus on entity-level or semantic block-level understanding, offering stronger abstraction capabilities while sacrificing some fine-grained detail. Striking a balance between detail and abstraction, joint-grained frameworks such as StructExt [structext], StructExtV2 [structextv2], and WuKong [wukong] have emerged, combining both granular and holistic representations. More recently, Large Language Model (LLM)-based architectures—including LayoutLLM [layoutllm] and HRVDA [hrvda]—have been proposed to support cross-modal reasoning and complex semantic understanding through instruction tuning and prompt-driven paradigms.
In this survey, we provide a comprehensive review of deep learning-based VRD-CU methodologies, tracing their evolution from early neural encoders to current multimodal transformers and LLM-based frameworks. We categorize existing approaches based on task formulation, model architecture, feature integration strategies, and pretraining objectives, comparing their effectiveness, generalisability, and scalability across benchmark datasets. We further highlight emerging trends, challenges, and future directions, aiming to support both researchers and practitioners in navigating this dynamic and impactful research area.
1.2 Scope
This survey aims to provide a comprehensive and critical review of recent advances in deep learning-based Visually Rich Document Content Understanding (VRD-CU). We focus on approaches that leverage neural architectures to understand and reason over the content of complex VRDs. The scope of this survey is defined by the following criteria:
  • Focus on Content-Level Understanding Tasks: We include works that address document content understanding tasks, specifically Key Information Extraction (KIE), Visually Rich Document Question Answering (VRD-QA), Entity Linking (EL). We exclude studies that are limited to document structure analysis (e.g., layout segmentation or reading order detection) without directly contributing to semantic content understanding.
  • Deep Learning-based Models: We primarily review papers that propose novel deep learning methods tailored for VRD-CU tasks. While Section 2.2.1 briefly summarises early heuristic and traditional machine learning approaches for historical context, the core focus remains on modern neural frameworks.
  • Dataset Coverage: We compile and analyse publicly available benchmark datasets for VRD-CU tasks, particularly those introduced or widely adopted in top-tier conferences and journals since 2019. These datasets support content-level reasoning in multimodal documents.
@@ 1.2 Scope
  • Emphasis on Multimodal and Multi-page Understanding: In contrast to previous surveys that focus narrowly on subtasks or single-page settings, this paper highlights models capable of multimodal reasoning and handling entire, multi-page documents holistically, better reflecting real-world document comprehension needs.
@@
This survey is designed to support researchers and practitioners in understanding the current landscape of deep learning models for VRD-CU, identifying emerging trends, and navigating the key resources, including models and datasets, which define this growing field. It should be noted that this article only investigates models and datasets designed for the entire document or the document page rather than the specific document components such as table, chart. Therefore, models and datasets proposed for table detection, table structure recognition, and chart or plot question answering will not be summarized in the main body of this survey.
1.3 Related Surveys
@@
Several surveys offer a comprehensive overview of general document understanding tasks. These surveys primarily focus on document layout analysis [subramani2020survey], table extraction [liu2023tabular], named entity recognition [ehrmann2023named], and document image analysis [lombardi2020deep] across diverse document types, including invoices [saout2024overview] and historical document [ehrmann2023named, lombardi2020deep]. In particular, computer vision-based research focuses on scanned document analysis and structure understanding. While these studies have advanced document image analysis, they often focus on fragmented subtasks and fall short of providing a holistic understanding of an entire document on multiple pages.
@@
Recent advancements in deep learning have fueled the emergence of VRD-CU tasks, which demand complex document content understanding capabilities. These tasks include key information extraction, question answering, and document entity linking. However, existing surveys have not adequately addressed the unique challenges and opportunities presented by VRD-CU, with a specific emphasis on deep learning-based multimodal approaches [cui2021document]. To bridge this gap, this survey aims to provide a comprehensive overview of VRD-CU frameworks and datasets, including multimodal feature extractions and fusions in both mono and multi-task VRD models.
1.4 Contributions
@@
  • The paper provides a detailed review and systematic categorization of VRD-CU frameworks and benchmark datasets, organized based on adopted strategies and downstream tasks.
  • It critically examines and compares different techniques used in VRD-CU models, focusing on feature representation and fusion, model architecture, and pretraining methods, highlighting their strengths, limitations, and appropriate scenarios.
@@
1.5 Survey Structure
Section 1 introduces the background of VRD-CU, outlining the aim, scope, and key contributions of this survey. Section 2 provides essential background knowledge, including definitions of the VRD-CU tasks, the evolution of document understanding techniques and terminology definition. Sections 3 and 3 review notable mono-task and multi-task frameworks, respectively. Section 5 presents benchmark datasets across three VRD-CU subtasks with evaluation metrics description. Section 6 offers a critical analysis of the reviewed models, highlighting their strengths and limitations through quantitative evaluation. Section LABEL:sec:challenges identifies emerging trends and key challenges, with a focus on practical applications. Finally, Section LABEL:sec:conclusion summarizes the survey findings.
2 Background
2.1 Task Definition
Based on their purpose and application scenarios, visually rich document content understanding (VRD-CU) tasks can be categorized into three main types: Key Information Extraction, Question Answering, and Entity Linking, as illustrated in Figure 1.
Figure 1: Visually rich document content understanding task clarifications.
@@ 2.1 Task Definition
  • Entity Linking: refers to identifying the semantic relations between document semantic entities to construct the logical structure of the input document image.
  • Question Answering: is a task of answering questions about a VRD by using natural languages. Based on the answer types, it can be divided into extractive QA and generative QA. The answer from extractive QA is directly extracted from the target document, while generative QA requires generating answers based on comprehensively understanding questions and related VRDs.
2.2 Development of Document Understanding
2.2.1 Traditional Approaches
@@
Rule-based methods, as highlighted in several studies [watanabe1995layout, o1993document, seki2007information, rusinol2013field], have demonstrated high precision in domain-specific applications. However, these methods have several drawbacks: they are manually intensive, costly, and require expert intervention for tailored customization. Additionally, they are inflexible, often necessitating frequent manual updates, even for minor modifications. In response to these limitations, machine learning-based approaches have been proposed for document understanding. For instance, SVM-based methods have been utilized for layout understanding [oliveira2017fast], and TF-IDF techniques combined with hand-crafted features have been applied for extracting information from invoices. Furthermore, rule-based and statistical models have been used for entity extraction. Despite these advancements, machine learning methods rely heavily on human intervention and domain-specific expertise. They are time-consuming and often deliver suboptimal performance. Moreover, the majority of these methods typically depend on single-modality data, restricting them to either layout, text, or visual inputs.
2.2.2 Single Modality-based Approaches
@@
With the advancement of deep learning, deeper model architectures such as CNNs [chargrid, yang2017learning] have emerged, and pretrained language [bert, roberta] and vision models [fasterrcnn, maskrcnn] are now commonly used as robust baselines for understanding VRD content and structure. Due to the multimodal nature of VRDs, which involves the integration of text, vision, and layout, researchers are increasingly focusing on leveraging this combined information to achieve significant improvements in various downstream VRD-CU tasks.
2.2.3 Cross Modality based Approaches
@@
Considering the multimodal nature of VRDs, many frameworks propose various methods to encode multimodal information, including text, vision, and layout, and fuse them effectively. Text and vision information is normally encoded by pretrained backbones such as BERT [bert] or RoBERTa [roberta] for textual features, Faster-RCNN [fasterrcnn] or Mask-RCNN [maskrcnn] for visual features. For layout information, different encoding methods are introduced, including linear projection [docstruct], 2D positional encoding [layoutlm], and attention bias to allow the proposed models to be layout-sensitive. Different feature fusion methods are introduced including summing up [pick], concatenation [formnet], attention-based contextual learning [majumder2020representation], and prompting [icld3ie]. However, most of those frameworks leverage implicit knowledge from pretrained backbones with a task-orientated shadow adapter for specific VRD-CU downstream tasks such as KIE [formnetv2, laser, chen2023task, genkie] or EL [zhang2021entity, kvpformer, carbonell2021named]. Those frameworks tend to achieve delicate performance on specific tasks or document formats instead of acquiring a generalised model to represent documents comprehensively.
2.2.4 Multimodal Pre-training Approches
@@
Inspired by the success of BERT-style models [bert, roberta] in acquiring knowledge through self-supervised learning, pretrained document understanding models have emerged to harness self-supervised or supervised pretraining tasks from extensive document collections. LayoutLM [layoutlm], the first encoder-only VRD-CU model, utilizes self-supervised tasks, such as masked vision-language modelling, with text and layout information. Subsequent models have expanded on this by integrating layout information [lilt, layoutmask, structext] and visual cues [layoutlmv2, layoutlmv3] through multimodal transformers. While encoder-only models have shown significant improvements on various benchmark datasets [funsd, cord, sroie, docvqa, rvlcdip], they often require detailed annotations and are limited by fixed input lengths. To address these limitations, encoder-decoder frameworks [udop, donut, dessurt] and prompt-based methods for LLMs/MLLMs [layoutllm, hrvda] have been developed, enhancing layout awareness and performance in VRD-CU tasks. However, a significant gap remains in effectively applying these models in real-world scenarios with zero shot.
2.3 Terminology Definition
After reviewing the evolution of VRD-CU, we introduce standardized terminology to categorize existing frameworks along three key dimensions: Downstream Task Focus, Model Architecture, Textual Granularity and Visual Information.
  • Downstream Task Focus: Frameworks are categorized into mono-task and multi-task types. Mono-task frameworks (see Section 3) are tailored for a single VRD-CU task (e.g., KIE or VQA) with task-specific designs, while multi-task frameworks (see Section 3) are designed to handle multiple tasks simultaneously, typically leveraging pretraining or shared components to improve generalizability.
  • Model Architecture: Frameworks can be grouped into encoder-only, encoder-decoder, and decoder-only categories. Encoder-only models focus on representation learning; encoder-decoder models generate task-specific outputs from document inputs; decoder-only models use autoregressive decoding for end-to-end generation.
  • Textual Granularity: Depending on how document structure is represented, frameworks process inputs at different levels: fine-grained (word-level), coarse-grained (e.g., entity-level inputs such as paragraphs, tables, or section headers), or joint-grained, which combines both. Fine-grained inputs capture detailed spatial and lexical information through individual word tokens, while coarse-grained representations emphasize higher-level semantic structures and contextual dependencies.
  • Visual Information: For frameworks that incorporate visual information, we categorize them based on the visual feature extraction methods: RoI-based and patch-based. RoI-based frameworks often use object detectors or document parsers to identify Regions of Interest (RoIs) corresponding to semantic entities, followed by a CNN (e.g., ResNet) to extract visual features via RoI-Align—commonly referred to as RoI-based frameworks. In contrast, Transformer-based frameworks (e.g., ViT [vit], Swin-Transformer [liu2021swin]) typically divide the document image into uniform patches and encode them sequentially, forming the basis of patch-based frameworks.
These terms provide a unified lens to understand and compare the design choices across VRD-CU frameworks in Section 3, Section 3 and Section 6.
3 Mono-Task Document Understanding Frameworks
VRD-CU comprises several distinct downstream tasks, each designed to meet specific application needs and user requirements. This section introduces methods tailored to individual tasks. As illustrated in Figure 2, we focus on three core VRD-CU tasks—Key Information Extraction (KIE), Entity Linking (EL), and Question Answering (QA)—and summarize corresponding models along with key trends and insights in their development.